A galvanometer of internal resistance 100Ω gives a full scale deflection for a current of 10 mA. Calculate the values of the resistance necessary to convert the galvanometer to an ammeter reading up to 10 A


1
Expert's answer
2021-10-06T08:15:08-0400

Given:

"R_0=100\:\Omega"

"I_0=10\:\rm mA=0.01\: A"

"I=10\:\rm A"


The resistance of the shunt is given by

"R=\frac{I_0}{I-I_0}R_0=\frac{0.01}{10-0.01}*100=0.1\:\Omega"